the plates of a parallel plate capacitor,5.0x10^-3m apart are maintained at a potential difference of 5.0x10^4 calculate the magnitude of the i. electric field intensity between the plates ii. force on the electron.
1
Expert's answer
2020-05-13T10:14:18-0400

Potential between plates is defined as "U=Ed" , where "U"-potential between plates, "E"-electric field, "d"-distance between plates. Thus, "E=\dfrac{U}{d}=\dfrac{5\cdot 10^4}{5\cdot 10^{-3}}=10^{7} ~V/m". Force on electron "F=qE\approx1.6\cdot10^{-19}\cdot10^7\approx1.6\cdot 10^{-12}"